You are given the following probability density function, 6x(r), for the cosine of the surface angle, S, of a laser etching tool. The distribution function has one parameter, a, and one constant, c. -1sxs1 a) What is the value of the constant, e? b) What is the moment estimator for a? c) Explain how you can determine if this moment estimator is unbiased. t... . 24) denote a random sample of sample size n 24 with sample mean of-0.01 and sample variance of 0.1, for values of x. Use the moment estimator to estimate α? e) Use the method of maximum likelihood to estimate the parameter a from a random sample. Just derive the equation, you do not have to find a closed form expression or value for a ) Which method would you use to estimate a? Explain your answer.
Problem #2 (5 points) The lifetimes of 6 components in an electronics device are known to be independent random variables governed by exponential distributions with means of 1000, 2000, 3000, 4000, 7000 and 10,000 hours, respectively. a) b) c) d) What is the probability that the lifetimes of all the components exceed 5000 hours? What are the mean and variance of life of all of the components? What is the probability that at least one component lifetime exceeds 10,000 hours? If the electronics device continues to operate only if all components operate, what is the mean life of the electronics device?
Answer #1


e ven Trobab ensi UnCtOn 貶(x)-.axc-1 ; ~1/121 2. C-24 UL IS 0 The Jecond moment ig30-5 5(a-31 fhe method oł moment for i 2 20 MoM + 5 lte moment Egtnoator i unbiased 0-24, y2 24 (ol tool 2) 2.4024leve 90(8 4024)/3 +5/ 2-5(2.4024) 1:933G5 OM = obden2 GComponer Asal! the Comporeok ave independaof ue Consider P (%; >5000 ,5000 1o000 -5000件000 0-114.2.0-0.5 0.29 Thus, the probailty that the lide times o allComponeots exceed s000 is 0.29 ⑥ nhe mea= 1-0, 00047 = 0.99953 Pobabiliy hat atleast One Componet lie tme eds 10,000 hour s alen t Une 0.99953 coidouice Coptioue to